How do I solve the following problem: The length of a rectangle is 2 cm less than 7 times the width. The perimeter is 60 cm. Find the width and length.
Answer by Theo(13342)

length is 2 centimeters less than the 7 times the width.
let l = length
let w = width
l = 7w - 2
perimeter = 60 cm = 2 * (length + width) = 2 * (l + w).
since l = 7w - 2, then perimeter = 2 * (7w - 2 + w).
combine like terms to get perimeter = 2 * (8w - 2).
simplify to get perimeter = 16w - 4.
since perimeter = 60 centimeters, then 60 = 16w - 4
add 4 to both sides of this equation to get 64 = 16w
solve for w to get w = 64 / 16 = 4.
since l = 7w - 2, then l = 7*4 - 2 = 28 - 2 = 26.
you have length = 26 and width = 4.
perimeters = 2 * (length + width) = 2 * (26 + 4) = 2 * 30 = 60.
solution is confirmed to be good.
width = 4
length = 26